# Unveiling the Learning Mind of Language Models: A Cognitive Framework and Empirical Study

### v1

@@ 6 Learning from Experience (LfE)
We investigate Learning from Experience (LfE), where models adapt by accumulating and utilizing prior interaction history. We evaluate this ability in two settings: (1) Experience-Driven Adaptation in Competitive Games, where agents condition on past multi-round play to adapt strategic behavior; and (2) In-context Examples as Off-policy Exploration Supervisions, where LLMs learn from ICL-style demonstrations viewed as episodic traces.
@@ 6.1 Experience-Driven Adaptation in Competitive Games
We adopt the same experimental setup as in Section 5.1, using TextArena (Guertler et al., 2025), a suite of competitive environments featuring symbolic rules and multi-turn dynamics. The results are shown in Figure 5. Here, "with experience" refers to a setting where, during the $k$-th game, the player selects three prior game experiences from rounds 0 to $k{-}1$ (or all available ones if fewer than three exist), which are then used as ICL examples, allowing the model to learn from past experiences. Implementation details are provided in Appendix A.5.
@@ 6.2 In-context Examples as Off-policy Exploration Supervisions
To investigate how accumulated experience influences model behavior, we treat in-context examples as trajectories encoding implicit supervision. This setup parallels off-policy learning: unlike competitive gameplay in Section 6.1, where models adapt through active interaction (on-policy), here they condition passively on externally provided experience examples. In Figure 6(a), we evaluate four base models, LLaMA3.1-8B, Qwen2.5-7B, Mistral-7B, and Phi-3-mini, under varying numbers of in-context exemplars. In Figure 6(b), we scale the amount of examples up to 1300 and track the performance of Qwen2.5-14B-Instruct-1M as well as Qwen2.5-7B-Instruct-1M variants. “Norm acc" in Figure 6(b) denotes accuracy normalized by the model’s zero-shot performance. Finally, in Figure 6(c), we compare two supervision regimes: in-context learning and instruction tuning. We use MagpieMath (Xu et al., 2024a), as introduced in Section 4, as the source of experience (i.e., the ICL examples), and the test set is also drawn from it. All ICL-example experiments were run 20 times, and we report the mean and standard deviation. Implementation details are provided in Appendix A.5.
@@
As shown in Figure 6(a), all models exhibit a non-monotonic trend as the number of in-context examples increases: performance improves initially, peaks, and then declines. This highlights that current LLMs are effective few-shot learners but struggle to scale with longer trajectories, raising doubts about their many-shot capabilities. Figure 6(b) further confirms this limitation: when the number of in-context examples exceeds  900, performance drops sharply, even in larger models. This suggests that excessive exampls may exceed attention capacity or interfere with generalization. Additionally, the 14B model consistently outperforms the 7B model, suggesting greater robustness to many-shot degradation. Figure 6(c) compares few-shot learning and instruction tuning. When data is limited, few-shot learning is highly efficient, using just 3 examples matches the performance of tuning with 5k instances. However, as data increases, instruction tuning continues to improve, while few-shot performance declines. This suggests a practical strategy: use few-shot learning in low-data regimes, and switch to tuning as more data becomes available.
@@ 7 LearnArena: A Benchmark Suite for General Learning Ability
Experiment Setup.
We construct a benchmark suite that evaluates LLMs’ general learning ability across three learning dimensions. The benchmark is built upon a modified version of the TextArena framework (Guertler et al., 2025), where each environment is cast as a two-player game between Player-0 and Player-1. We fix Player-0 as Qwen2.5-32B and designate the evaluated LLM as Player-1. At each round $k$, Player-1 selects three experiences from the previous $k-1$ games and submits them to Player-0 for feedback. Player-0 provides suggestions for improvement, which Player-1 incorporates in the current round, representing Learning from Instructor. Separately, a concise summary of the game rules is generated by Qwen2.5-32B and given to Player-1, capturing Learning from Concept. In addition, Player-1 is encouraged to perform its own analysis of the selected past experiences and apply its conclusions in the current game, reflecting Learning from Experience. We evaluate each model-environment pair across 20 independent matches and report the average win rate of Player-1. The environments span a diverse set of strategic and social reasoning tasks, including Checkers (CH), Stratego (ST), Tic Tac Toe (TT), Truth and Deception (TD), SpellingBee (SB), SpiteAndMalice (SM), Tak (TK), and WordChains (WC). We evaluate the following models: Llama-3.1-8B, Mistral-7B-v0.3, Mistral-8B-2410, Qwen2.5-7B, Qwen2.5-14B-Instruct, Qwen2.5-32B-Instruct, Qwen3-8B, Qwen3-14B, GPT-4-o, and GPT-4-o-mini. Further implementation details are provided in Appendix A.6.

### v3

@@ 6 Learning from Experience (LfE)
We investigate Learning from Experience (LfE), where models adapt by accumulating and utilizing prior interaction history. We evaluate this ability in two settings: (1) Experience-Driven Adaptation in Competitive Games, where agents condition on past multi-round play to adapt strategic behavior; and (2) In-context Examples as Off-policy Exploration Supervision, where LLMs learn from ICL-style demonstrations viewed as episodic traces.
@@ 6.1 Experience-Driven Adaptation in Competitive Games
We adopt the same experimental setup as in Section 5.1, using TextArena (guertler2025textarena), a suite of competitive environments featuring symbolic rules and multi-turn dynamics. The results are shown in Figure 6. Here, "with experience" refers to a setting where, during the $k$-th game, the player selects three prior game experiences from rounds 0 to k1k{-}1 (or all available ones if fewer than three exist), which are then used as ICL examples, allowing the model to learn from past experiences. Implementation details are provided in Appendix A.5.
@@ 6.2 In-context Examples as Off-policy Exploration Supervisions
To investigate how accumulated experience influences model behavior, we treat in-context examples as trajectories encoding implicit supervision. This setup parallels off-policy learning: unlike competitive gameplay in Section 6.1, where models adapt through active interaction (on-policy), here they condition passively on externally provided experience examples. In Figure 7(a), we evaluate four base models, LLaMA3.1-8B, Qwen2.5-7B, Mistral-7B, and Phi-3-mini, under varying numbers of in-context exemplars. In Figure 7(b), we scale the amount of examples up to 1300 and track the performance of Qwen2.5-14B-Instruct-1M as well as Qwen2.5-7B-Instruct-1M variants. “Norm acc" in Figure 7(b) denotes accuracy normalized by the model’s zero-shot performance. Finally, in Figure 7(c), we compare two supervision regimes: in-context learning and instruction tuning. We use MagpieMath (xu2024magpie), as introduced in Section 4, as the source of experience (i.e., the ICL examples), and the test set is also drawn from it. All ICL-example experiments were run 20 times, and we report the mean and standard deviation. Implementation details are provided in Appendix A.5.
@@
As shown in Figure 7(a), all models exhibit a non-monotonic trend as the number of in-context examples increases: performance improves initially, peaks, and then declines. This highlights that current LLMs are effective few-shot learners but struggle to scale with longer trajectories, raising doubts about their many-shot capabilities. Figure 7(b) further confirms this limitation: when the number of in-context examples exceeds  900, performance drops sharply, even in larger models. This suggests that excessive examples may exceed attention capacity or interfere with generalization. Additionally, the 14B model consistently outperforms the 7B model, suggesting greater robustness to many-shot degradation. Figure 7(c) compares few-shot learning and instruction tuning. When data is limited, few-shot learning is highly efficient, using just 3 examples matches the performance of tuning with 5k instances. However, as data increases, instruction tuning continues to improve, while few-shot performance declines. This suggests a practical strategy: use few-shot learning in low-data regimes, and switch to tuning as more data becomes available.
@@ 7 LearnArena: A Benchmark Suite for General Learning Ability
Experiment Setup.
We construct a benchmark suite that evaluates LLMs’ general learning ability across three learning dimensions. The benchmark is built upon a modified version of the TextArena framework (guertler2025textarena), where each environment is cast as a two-player game between Player-0 and Player-1. We fix Player-0 as Qwen2.5-32B and designate the evaluated LLM as Player-1. At each round $k$, Player-1 selects three experiences from the previous k1k-1 games and submits them to Player-0 for feedback. Player-0 provides suggestions for improvement, which Player-1 incorporates in the current round, representing Learning from Instructor. Separately, a concise summary of the game rules is generated by Qwen2.5-32B and given to Player-1, capturing Learning from Concept. In addition, Player-1 is encouraged to perform its own analysis of the selected past experiences and apply its conclusions in the current game, reflecting Learning from Experience. We evaluate each model-environment pair across 20 independent matches and report the average win rate of Player-1. The environments span a diverse set of strategic and social reasoning tasks, including Checkers (CH), Stratego (ST), Tic Tac Toe (TT), Truth and Deception (TD), SpellingBee (SB), SpiteAndMalice (SM), Tak (TK), and WordChains (WC). We evaluate the following models: Llama-3.1-8B, Mistral-7B-v0.3, Mistral-8B-2410, Qwen2.5-7B, Qwen2.5-14B-Instruct, Qwen2.5-32B-Instruct, Qwen3-8B, Qwen3-14B, GPT-4-o, and GPT-4-o-mini. Further implementation details are provided in Appendix A.6.